OpenStudy (nene101):

A cruise boat travels 72 miles downstream in 3 hours and returns to its starting point upstream in 12 hours find speed of the stream ?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (nene101):

Thanks

likeabossssssss (likeabossssssss):

np what do u think it is ?

likeabossssssss (likeabossssssss):

@Idkanything u want to help him/her?

OpenStudy (idkanything):

@likeabossssssss ya

likeabossssssss (likeabossssssss):

ok do it just give direct answer ok

OpenStudy (idkanything):

ok

OpenStudy (idkanything):

@3mar

OpenStudy (idkanything):

36 miles

OpenStudy (radar):

You are wanting the speed of the stream (current). Some facts that you must consider the trip is the same distance each way. Let x = speed of the stream. Let y = the speed of the boat (relative to the water) The combined speed when going down stream is 72/3 from speed equals distance divided by time. The speed was 24 mph downstream. The speed up stream was 72/12 or 6 mph. with these facts we can write our equations: x + y = 24 and y - x = 6 or -x + y = 6 add the equations eliminating x and solving for y 2y = 30, or y = 15 The speed of the boat is 15 mph. Now subtract the two equations eliminating y and solving for x. 2x = 18 or x = 9, the current is 9 mph.

OpenStudy (radar):

The speed of the stream is x or 9 mph